OpenStudy (anonymous):

Given f(x) = 2x2 + 2x and g(x) = x + 1, find the function (f/g)(x). a. 2x; excluded value: x = -1 b. 2x; excluded value: x = 0 c. 1/2x; excluded value: x = -1 d. 1/2x; excluded value: x = 0

OpenStudy (anonymous):

(f/g)(x) = f(x) / g(x) = (2x^2+2x)/(x+1) = (2x(x+1))/(x+1) = 2xexcluded value, x=-1

OpenStudy (anonymous):

thanks :)